a chemist has two solutions of h2so4. one has 40% concentration and the other has a 25% concentration. how many liters of solution must be mixed to obtain 96 liters of a 27% solution
Answer by Maths68(1474)

Soluton A
Amount = x
Concentration = 40%=0.4
================================================================================
Solution B
Amount = 96-x liters
Concentration = 25%=0.25
================================================================================
Resultant Solution
Amount=96
Concentration=27%=0.27
================================================================================
(Amount of A * Concentration of A)+(Amount of B * Concentration of B) = (Amount of Resultant * Concentration of Resultant)
(x*0.4)+(96-x)(0.25)=(96)*0.27
0.4x+24-0.25x=25.92
0.15x=25.92-24
0.15x=1.92
0.15x/0.65=1.92/0.15
x=12.8 liters

================================================================================
Soluton A
Amount = x = 12.8 liters
Concentration = 40%=0.4
================================================================================
Solution B
Amount = 96-x=96-12.8=83.2 liters
Concentration = 25%=0.25
===============================================================================
12.8 liters of 40% concentration and 83.2 liters of 25% concentration solutoins must be mixed to obtain 96 liters of a 27% solution